OpenStudy (anonymous):

Find the thirtieth term of the following sequence. -6, -4, -2, 0, ... 52 64 66

OpenStudy (welshfella):

this is an arithmetic series with first term -6 and common difference 2. This means that 2 is added to each term to get the next the formula for the nth term is a + d(n - 1) where a = first term , d = common difference and n = term number so if you plug in a = -6 , d = 2 and n = 30 into this formula you can work out thr 3oth term

OpenStudy (anonymous):

it would have been 60 right but the answers dont say 60

OpenStudy (welshfella):

its not 60 you need to calculate -6 + 2(30-1) - work out the subtraction in the parentheses first then multiply by 2 then add -6

OpenStudy (welshfella):

add ing -6 is the same as subtracting 6

OpenStudy (anonymous):

if its thirty more terms it would be either -60 or 60 depending on which direction you go in

OpenStudy (anonymous):

unless it means start at -6 and keep going.?

OpenStudy (welshfella):

just work out the expression i just wrote

OpenStudy (anonymous):

but at the end of the 0 theres ... like it wants me to go in that direction

OpenStudy (welshfella):

what is 30 - 1?

OpenStudy (anonymous):

its 52 but why?

OpenStudy (welshfella):

-6 + 2(29) = 1-6 + 58 = 52

OpenStudy (welshfella):

* -6 + 58 = 52

OpenStudy (anonymous):

ok nvrmind thx anyway i still got it right

OpenStudy (welshfella):

yw you can see that it will get up to high positive values pretty quickly -6 -4 -2 0 2 4 6 8 10 12 14....